OpenStudy (anonymous):

Quadrilateral OPQR is inscribed inside a circle as shown below. What equation would be needed to solve for angle P? What is the measure of angle P?

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@random231

random231 (random231):

um its showing error for meh!

OpenStudy (anonymous):

i tried to put the picture but it would not let me

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

A carnival ride is in the shape of a wheel with a radius of 30 feet. The wheel has 30 cars attached to the center of the wheel. What is the central angle, arc length, and area of a sector between any two cars?

OpenStudy (anonymous):

it don't have a picture i am trying to figure the other one out

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

for the quadrilateral problem, have a look at this http://www.geom.uiuc.edu/~dwiggins/pict47.GIF

jimthompson5910 (jim_thompson5910):

basically, if there's a quadrilateral inscribed in a circle, then the opposite angles add to 180 degrees

OpenStudy (ranga):

Just follow up on jim_thompson's suggestion. When a quadrilateral is inscribed in a circle, the opposite angles add to 180 degrees. Therefore, angle P + angle R = 180 degrees. From the diagram, angle P = y and angle R = 3y + 8 Therefore, y + 3y + 8 = 180 Solve for y and that will be angle P.

OpenStudy (anonymous):

How do you find y

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ranga):

y + 3y + 8 = 180 4y + 8 = 180 Subtract 8 from both sides. Then divide both sides by 4 to find y.